OpenStudy (briannabug):

The surface area of a cone with radius r units and slant height s units is shown below: Surface area = 3.14(rs + r2) Part A: If r = 6 units and s = 4 units, write an expression that can be used to calculate the surface area of the cone. Part B: What is the surface area of the cone? I don't understand this good, so any help would be great! Thanks!

jhonyy9 (jhonyy9):

are you here please ?

jhonyy9 (jhonyy9):

surface area SA = 3,14(rs +r^2) = 3,14(6*4 +6^2) = ?

OpenStudy (shelbs.13):

Plug in the numbers first then solve the equations

OpenStudy (briannabug):

Thanks!